OpenStudy (olivia.0498):

1/10 as a percentage and decimal

OpenStudy (olivia.0498):

help me

OpenStudy (steve816):

10% 0.10

OpenStudy (olivia.0498):

1/2 divided by 2/3

OpenStudy (olivia.0498):

help

OpenStudy (shadowlegendx):

\[\frac{ 1 }{ 2 } \div \frac{ 2 }{ 3 } = \frac{ 1 }{ 2 } \times \frac{ 3 }{ 2 }\]

OpenStudy (olivia.0498):

thanks.. prime factors for 35

OpenStudy (shadowlegendx):

\[\frac{ 3 }{ 4 }\] would be your final answer, unless you need it in decimal form

OpenStudy (sooobored):

35= 5x7 5 is prime, 7 is prime, no further factors end

OpenStudy (mathmale):

Olivia: Only ONE question per post, please; if you share several questions at once, it can get messy (and hard to follow the discussion).